# ClipTTA: Robust Contrastive Vision-Language Test-Time Adaptation

### v1

@@ 1 Introduction
Entropy minimization is the most common TTA objective tent ; eta ; sar ; rotta , as it mirrors the cross-entropy training of standard classifiers. However, it is fundamentally misaligned with the contrastive image-text pre-training objective of VLMs like CLIP, as illustrated in Fig. 1, potentially hindering adaptation due to differing gradient dynamics.  Recent works have attempted to improve TTA of CLIP by leveraging visual-textual similarities in a transductive manner clipartt ; watt , yet the objective misalignment remains unresolved. Furthermore, when labeled data is available, recent work on fine-tuning flyp demonstrates that using the exact same loss function as during CLIP pre-training leads to better performance on downstream tasks.
@@
Together, these observations raise a central question: how to design an adaptation loss that is more suited for gradient-based TTA of CLIP?
In this work, we introduce ClipTTA, a new test-time adaptation method tailored to vision-language models. It employs a soft contrastive image-text loss that mirrors CLIP’s pre-training objective, providing natural continuity in adaptation. As illustrated in Fig. 1, this design reflects our central assumption: adaptation losses should align with the model’s multimodal contrastive training paradigm. Importantly, the contrastive nature of the ClipTTA loss links predictions within a batch, incorporating mechanisms to mitigate the risk of class collapse caused by noisy pseudo-labels. It also demonstrates increased robustness in open-set scenarios, where both in-distribution (ID) and out-of-distribution (OOD) samples are present. We further augment it with a discriminative loss to separate ID from OOD samples, improving performance under open-set conditions.
@@ 3 ClipTTA
Figure 2: Illustration of ClipTTA. ClipTTA in Sec. 3.1 consists of a soft contrastive loss specifically designed for TTA of VLMs like CLIP. We show in Sec. 3.2 that ClipTTA is robust to class collapse and pseudo-label errors. Finally, we add an OCE loss to be robust to OOD samples in batches in Sec. 3.3 and improve the ID/OOD detection and accuracy in open-set scenarios.
We introduce ClipTTA, a contrastive test-time adaptation method tailored to VLMs such as CLIP, as illustrated in Fig. 2. By aligning the adaptation objective with CLIP’s image-text contrastive pre-training described in Sec. 3.1, ClipTTA improves robustness to pseudo-label errors and class collapse through its batch-aware formulation, as demonstrated by our gradient analysis in Sec. 3.2. Combined with the Outlier Contrastive Exposure loss introduced in Sec. 3.3, it improves both OOD detection and accuracy for robust adaptation in open-set scenarios.
@@ 3.1 Contrastive adaptation loss at test-time
Although Eq. 1 and Eq. 2 appear similar, they differ in their softmax normalization: Eq. 1 normalizes over $C$ classes, while Eq. 2 normalizes over the $N$ predicted classes in the batch.
@@
This loss retains CLIP’s contrastive structure while explicitly modeling uncertainty in pseudo-labels. As shown in  Fig. 2, the first term computes the entropy over the image-to-text probability distribution (row-wise), and the second term the entropy over the text-to-image probability distribution (column-wise) within the batch. Analogous to entropy minimization, which replaces hard cross-entropy with a soft and uncertainty-aware loss, our soft contrastive loss is a principled extension of the VLMs’ contrastive scheme. Furthermore, it demonstrates enhanced robustness to pseudo-label errors, as studied in Sec. 3.2. To ensure fair comparisons, we use only the image-to-text term of Eq. (3) in the main experiments, as most gradient-based TTA methods update only the visual encoder. The effect of simultaneously updating the text encoder is evaluated in Appendix C.
@@
Following prior TTA research SWR ; ostta ; stamp ; rotta , we also incorporate standard techniques such as entropy regularization and a class-wise confident memory (CCM) to enhance adaptation. The regularization loss, based on negative marginal entropy, diversifies the predictions by uniformizing the prediction distribution across classes. Defining $\bar{q}_{c}=\frac{1}{N}\sum_{i=1}^{N}q({\bm{t}}_{c}|{\bm{x}}_{i})$ as the batch-wise average probability for class $c$ (i.e., over probabilities in Eq. 1), the regularization loss is $\mathcal{L}_{\text{reg}}(\theta)=\sum_{c=1}^{C}\bar{q}_{c}\log\bar{q}_{c}$.  The final ClipTTA loss integrates the soft-contrastive loss Eq. (3), the regularization term, and the CCM memory.  Memory batches $\mathcal{M}$, equal in size to test batches, are used to compute the adaptation loss:
[…]
@@ 3.2 Gradient Analysis
Let $N_{k}$ be the number of samples in the batch pseudo-labeled as class $k$, and $q_{ik}=q({\bm{t}}_{k}|{\bm{x}}_{i})$ as in Eq. 1. The gradient of $\mathcal{L}_{\text{s-cont}}$ w.r.t. ${\bm{z}}_{v}^{i}$ is:
@@
See Sec. A.1. ∎
This expression shows that the gradient for sample ${\bm{x}}_{i}$ aggregates contributions from all pseudo-captions in the batch, each weighted by $\beta_{i,j}$. Each contribution consists of two effects. The first term, $-\widehat{{\bm{z}}_{t}}^{j}$, acts as an attractive force pulling ${\bm{z}}_{v}^{i}$ toward pseudo-caption $\hat{{\bm{t}}}_{j}$. In contrast, the second term, $\sum_{k}w_{k,i}{\bm{z}}_{t}^{k}$, introduces a repulsive force that pushes the embedding away from dominant class directions since classes that are more frequently predicted exert stronger repulsion.
Importantly, the coefficients $\beta_{i,j}$ are key to allowing the gradient of a sample to point toward a class different from its pseudo-label, enabling error correction by leveraging predictions from other samples in the batch, as illustrated on a toy dataset in Appendix B. They amplify the contribution of confident and semantically similar pairs in the gradient update, allowing the model to rely on more reliable examples. For instance, if ${\bm{x}}_{i}$ is misclassified as class $k^{\prime}$ but is close to another sample ${\bm{x}}_{j}$ whose pseudo-caption reflects the correct class $k$, a large $\beta_{i,j}$ steers the update toward ${\bm{z}}_{t}^{k}$. Such correction is not achievable by sample-wise objectives like TENT, which systematically reinforce the predicted class regardless of its correctness.
@@
$||\nabla_{{\bm{z}}_{v}^{i}}\mathcal{L}_{\text{s-cont}}||\underset{N_{k}\rightarrow N}{\rightarrow}0.$ (6)
@@
See Sec. A.1. ∎
To further illustrate, consider a binary classification setting with classes $a$ and $b$, where $a$ is the most predicted class in the batch (i.e., $N_{a}\gg N_{b}$). In that case, the gradient in Eq. 5 becomes:
$\nabla_{{\bm{z}}_{v}^{i}}\mathcal{L}_{\text{s-cont}}=[\beta_{i,a}q_{ib}-\beta_{i,b}q_{ia}]\frac{N_{a}N_{b}}{N_{a}q_{ia}+N_{b}q_{ib}}({\bm{z}}_{t}^{b}-{\bm{z}}_{t}^{a}).\$ (7)
@@
This self-regulation property acts as a built-in dampening mechanism that slows adaptation before collapse occurs, helping prevent convergence to dominant classes, preserving stable updates, and giving the model a chance to recover from poor pseudo-labeling. In contrast, entropy-based objectives such as TENT continue to reinforce dominant class predictions even as imbalance increases, accelerating collapse rather than preventing it (see derivation in Sec. A.1).
@@ 3.3 Outlier Contrastive Exposure loss
[…]
@@ 4 Experiments
ClipTTA is evaluated on four families of adaptation benchmarks: corruptions (CIFAR-10/100-C, Imagenet-C) with 15 perturbations, domain shifts (VisDA-C, PACS, OfficeHome, Imagenet-Domains), semantic datasets, including coarse- (CIFAR-10/100) and fine-grained classification (Imagenet, and 10 datasets from the CLIP zero-shot suite). In total, this represents a thorough evaluation over 75 datasets. A detailed description is provided in Sec. C.2. In open-set TTA, SVHN and Places-365 serve as OOD counterparts for CIFAR-10/100 and Imagenet, respectively.
@@
We use ViT-B/16 as CLIP’s backbone in all experiments. Adaptation is performed with batches of 128 images using the Adam optimizer and a learning rate of $10^{-4}$ over 10 iterations. Experiments are conducted in a non-episodic manner, i.e., without restoring the model’s parameters after each batch. Following the standard TTA protocol, we adapt the affine parameters of the visual encoder’s normalization layers. In the open-set setting, we add 128 OOD images per batch, as done in prior work unient ; stamp . The regularization and OCE losses’ weights are set to $\lambda_{reg}=1$ and $\lambda_{oce}=1$, respectively. We validate that ClipTTA is stable to variations of its hyper-parameters in Appendix C. Experiments were performed on two NVIDIA V100 32GB GPUs.
@@ 4.1 Main results
We compare ClipTTA with the two prevailing families of test-time objectives: (i) TENT-style losses, including TENT tent , ETA eta , SAR sar and RoTTA rotta , and (ii) CLIPArTT-derived losses, such as CLIPArTT clipartt and WATT watt . Table 1 reports top-1 accuracy under synthetic corruptions and real domain shifts. We highlight two key findings. First, ClipTTA substantially improves over zero-shot CLIP, with large gains when initial accuracy is low: +20.5 pts on CIFAR-10-C, +17.4 pts on CIFAR-100-C, and +15.6 pts on Imagenet-C. In contrast, other methods perform poorly under the same conditions, which can be attributed to the increased likelihood of class collapse and pseudo-label drift when initial accuracy is low. Further analysis and finer-grained experimental evidence are presented in Sec. A.1 and Appendix B, respectively. Second, ClipTTA is the only method that consistently achieves top performance across all benchmarks. While competing methods demonstrate strengths in specific scenarios, they fall short overall. For example, ETA performs best among the TENT-style methods on domain-shift datasets but still lags by 1.4 pts on average. Similarly, CLIPArTT is most competitive on corruption benchmarks but remains 8.3 pts behind. This trend persists across both coarse- and fine-grained datasets (Fig. 3, with extended results in Sec. C.3). Altogether, these results establish our soft contrastive loss as the most reliable and broadly effective objective for gradient-based TTA of CLIP.
@@
Figure 3: TTA results on semantic datasets. Top-1 accuracy on coarse-grained datasets (CIFAR-10 and CIFAR-100) (a) and on 11 fine-grained datasets (including Imagenet) (b). Comparison with gradient-based TTA methods tent ; eta ; sar ; rotta ; clipartt ; watt and alternative state-of-the-art TTA methods tpt ; tda .
@@
We further benchmark ClipTTA against two recent state-of-the-art TTA methods tailored to CLIP: TPT tpt , which adapts through text prompt tuning instead of updating normalization parameters, and TDA tda , a gradient-free approach that adjusts CLIP’s logits using cached predictions. As shown in Table 2, ClipTTA improves top-1 accuracy by an average of +19.4 pts over TPT and +15.6 pts over TDA. It achieves state-of-the-art results on nearly all domain-shift benchmarks, with the sole exception of the ImageNet-D suite, where TDA benefits from per-dataset hyperparameter tuning. However, we note that TDA lags far behind on corruption datasets, a standard benchmark in TTA, with an average gap of over 15 pts compared to ClipTTA. Figure 3 further confirms CLIPTTA’s advantage across both coarse- and fine-grained recognition tasks. While TPT and TDA are designed for single-image batches, our analysis in Sec. C.3 shows that ClipTTA remains competitive even in this challenging setting, thanks to its use of the CCM memory and maintains stable performance across a wide range of batch sizes.
@@ C-10-C C-100-C Imagenet-C Average VisDA-C PACS OfficeHome Imagenet-D Average
Table 3 presents results on the open-set scenario on Imagenet, where OOD detection needs to be performed alongside classification. First, we note that all closed-set methods, except ours, perform noticeably worse than zero-shot CLIP in OOD detection, highlighting CLIPTTA’s strong robustness to OOD sample contamination during adaptation. Notably, TENT and CLIPArTT suffer severe performance degradation in both classification and OOD detection, likely due to outlier interference in their pseudo-labeling process. Second, when equipped with our OCE loss, ClipTTA consistently outperforms specialized open-set TTA methods, which use heuristic OOD detection mechanisms, achieving +2.3 points AUC over UniEnt and +8.4 points AUC over SoTTA. Our soft contrastive objective reliably preserves and improves both accuracy and OOD detection, unlike these entropy-based methods, which tend to degrade CLIP’s initial performance. Additional results on other datasets are reported in Sec. C.3.

### v2

@@ 1 Introduction
Entropy minimization is the most common TTA objective tent ; eta ; sar ; rotta , as it mirrors the cross-entropy training of standard classifiers. However, it is fundamentally misaligned with the contrastive image-text pre-training objective of VLMs like CLIP, as illustrated in Fig.˜1, potentially hindering adaptation due to differing gradient dynamics.  Recent works have attempted to improve TTA of CLIP by leveraging visual-textual similarities in a transductive manner clipartt ; watt , yet the objective misalignment remains unresolved. Furthermore, when labeled data is available, recent work on fine-tuning flyp demonstrates that using the exact same loss function as during CLIP pre-training leads to better performance on downstream tasks.
@@
Together, these observations raise a central question:
How to design an adaptation loss that is more suited for gradient-based test-time adaptation of contrastive vision-language models such as CLIP?
In this work, we introduce ClipTTA, a new test-time adaptation method tailored to vision-language models. It employs a soft contrastive image-text loss that mirrors CLIP’s pre-training objective, providing natural continuity in adaptation. As illustrated in Fig.˜1, this design reflects our central assumption: adaptation losses should align with the model’s multimodal contrastive training paradigm. Importantly, the contrastive nature of the ClipTTA loss links predictions within a batch, incorporating mechanisms to mitigate the risk of class collapse caused by noisy pseudo-labels. It also demonstrates increased robustness in open-set scenarios, where both in-distribution (ID) and out-of-distribution (OOD) samples are present. We further augment it with a discriminative loss to separate ID from OOD samples, improving performance under open-set conditions.
@@ 3 ClipTTA
Figure 2: Illustration of ClipTTA. ClipTTA in Sec.˜3.1 consists of a soft contrastive loss specifically designed for TTA of VLMs like CLIP. We show in Sec.˜3.2 that ClipTTA is robust to class collapse and pseudo-label errors. Finally, we add an OCE loss to be robust to OOD samples in batches in Sec.˜3.3 and improve the ID/OOD detection and accuracy in open-set scenarios.
We introduce ClipTTA, a contrastive test-time adaptation method tailored to VLMs such as CLIP, as illustrated in Fig.˜2. By aligning the adaptation objective with CLIP’s image-text contrastive pre-training described in Sec.˜3.1, ClipTTA improves robustness to pseudo-label errors and class collapse through its batch-aware formulation, as demonstrated by our gradient analysis in Sec.˜3.2. Combined with the Outlier Contrastive Exposure loss introduced in Sec.˜3.3, it improves both OOD detection and accuracy for robust adaptation in open-set scenarios.
@@ 3.1 Contrastive adaptation loss at test-time
Although Eq.˜1 and Eq.˜2 appear similar, they differ in their softmax normalization: Eq.˜1 normalizes over $C$ classes, while Eq.˜2 normalizes over the $N$ predicted classes in the batch.
@@
This loss retains CLIP’s contrastive structure while explicitly modeling uncertainty in pseudo-labels. As shown in  Fig.˜2, the first term computes the entropy over the image-to-text probability distribution (row-wise), and the second term the entropy over the text-to-image probability distribution (column-wise) within the batch. Analogous to entropy minimization, which replaces hard cross-entropy with a soft and uncertainty-aware loss, our soft contrastive loss is a principled extension of the VLMs’ contrastive scheme. Furthermore, it demonstrates enhanced robustness to pseudo-label errors, as studied in Sec.˜3.2. To ensure fair comparisons, we use only the image-to-text term of Eq. (3) in the main experiments, as most gradient-based TTA methods update only the visual encoder. The effect of simultaneously updating the text encoder is evaluated in Appendix˜C.
@@
Following prior TTA research SWR ; ostta ; stamp ; rotta , we also incorporate standard techniques such as entropy regularization and a class-wise confident memory (CCM) to enhance adaptation. The regularization loss, based on negative marginal entropy, diversifies the predictions by uniformizing the prediction distribution across classes. Defining $\bar{q}_{c}=\frac{1}{N}\sum_{i=1}^{N}q({\bm{t}}_{c}|{\bm{x}}_{i})$ as the batch-wise average probability for class $c$ (i.e., over probabilities in Eq.˜1), the regularization loss is $\mathcal{L}_{\text{reg}}(\theta)=\sum_{c=1}^{C}\bar{q}_{c}\log\bar{q}_{c}$.  The final ClipTTA loss integrates the soft-contrastive loss Eq. (3), the regularization term, and the CCM memory.  Memory batches $\mathcal{M}$, equal in size to test batches, are used to compute the adaptation loss:
$\mathcal{L}_{\text{C{lip}TTA}}(\theta)=\frac{1}{2}\Big[\mathcal{L}_{\text{s-cont}}(\theta)+\mathcal{L}_{\text{s-cont}}^{\mathcal{M}}(\theta)\Big]+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}(\theta),$ (4)
@@ 3.2 Gradient Analysis
Let $N_{k}$ be the number of samples in the batch pseudo-labeled as class $k$, and $q_{ik}=q({\bm{t}}_{k}|{\bm{x}}_{i})$ as in Eq.˜1. The gradient of $\mathcal{L}_{\text{s-cont}}$ w.r.t. ${\bm{z}}_{v}^{i}$ is:
@@
See Sec.˜A.1. ∎
This expression shows that the gradient for sample ${\bm{x}}_{i}$ aggregates contributions from all pseudo-captions in the batch, each weighted by $\beta_{i,j}$. Each contribution consists of two effects. The first term, $\widehat{{\bm{z}}_{t}}^{j}$, acts as an attractive force pulling ${\bm{z}}_{v}^{i}$ toward pseudo-caption $\hat{{\bm{t}}}_{j}$. In contrast, the second term, $\sum_{k}w_{k,i}{\bm{z}}_{t}^{k}$, introduces a repulsive force that pushes the embedding away from dominant class directions since classes that are more frequently predicted exert stronger repulsion.
Importantly, the coefficients $\beta_{i,j}$ are key to allowing the gradient of a sample to point toward a class different from its pseudo-label, enabling error correction by leveraging predictions from other samples in the batch, as illustrated on a toy dataset in Appendix˜B. They amplify the contribution of confident and semantically similar pairs in the gradient update, allowing the model to rely on more reliable examples. For instance, if ${\bm{x}}_{i}$ is misclassified as class $k^{\prime}$ but is close to another sample ${\bm{x}}_{j}$ whose pseudo-caption reflects the correct class $k$, a large $\beta_{i,j}$ steers the update toward ${\bm{z}}_{t}^{k}$. Such correction is not achievable by sample-wise objectives like TENT, which systematically reinforce the predicted class regardless of its correctness.
@@
[…]
@@
See Sec.˜A.1. ∎
To further illustrate, consider a binary classification setting with classes $a$ and $b$, where $a$ is the most predicted class in the batch (i.e., $N_{a}\gg N_{b}$). In that case, the gradient in Eq.˜5 becomes:
$\nabla_{{\bm{z}}_{v}^{i}}\mathcal{L}_{\text{s-cont}}=[\beta_{i,a}q_{ib}-\beta_{i,b}q_{ia}]\frac{N_{a}N_{b}}{N_{a}q_{ia}+N_{b}q_{ib}}({\bm{z}}_{t}^{b}-{\bm{z}}_{t}^{a}).\\$ (7)
@@
This self-regulation property acts as a built-in dampening mechanism that slows adaptation before collapse occurs, helping prevent convergence to dominant classes, preserving stable updates, and giving the model a chance to recover from poor pseudo-labeling. In contrast, entropy-based objectives such as TENT continue to reinforce dominant class predictions even as imbalance increases, accelerating collapse rather than preventing it (see derivation in Sec.˜A.1).
@@ 3.3 Outlier Contrastive Exposure loss
In the open-set scenario, our optimization objective then becomes $\min_{\theta,\alpha}\mathcal{L_{\text{C{lip}TTA}}}+\lambda_{\text{oce}}\mathcal{L}_{\text{OCE}}$, where we update the parameters of the model $\theta$ and the ID / OOD threshold parameter $\alpha$ in an end-to-end fashion. Our OCE loss differs from the UniEnt loss unient since it is purely discriminative, enforcing a more direct separation between ID and OOD features, and since it learns the separation threshold $\alpha$.
@@ 4 Experiments
ClipTTA is evaluated on four families of adaptation benchmarks: corruptions (CIFAR-10/100-C, Imagenet-C) with 15 perturbations, domain shifts (VisDA-C, PACS, OfficeHome, Imagenet-Domains), semantic datasets, including coarse- (CIFAR-10/100) and fine-grained classification (Imagenet, and 10 datasets from the CLIP zero-shot suite). In total, this represents a thorough evaluation over 75 datasets. A detailed description is provided in Sec.˜C.2. In open-set TTA, SVHN and Places-365 serve as OOD counterparts for CIFAR-10/100 and Imagenet, respectively.
@@
We use ViT-B/16 as CLIP’s backbone in all experiments. Adaptation is performed with batches of 128 images using the Adam optimizer and a learning rate of $10^{-4}$ over 10 iterations. Experiments are conducted in a non-episodic manner, i.e., without restoring the model’s parameters after each batch. Following the standard TTA protocol, we adapt the affine parameters of the visual encoder’s normalization layers. In the open-set setting, we add 128 OOD images per batch, as done in prior work unient ; stamp . The regularization and OCE losses’ weights are set to $\lambda_{reg}=1$ and $\lambda_{oce}=1$, respectively. We validate that ClipTTA is stable to variations of its hyper-parameters in Appendix˜C. Experiments were performed on two NVIDIA V100 32GB GPUs.
@@ 4.1 Main results
We compare ClipTTA with the two prevailing families of test-time objectives: (i) TENT-style losses, including TENT tent , ETA eta , SAR sar and RoTTA rotta , and (ii) CLIPArTT-derived losses, such as CLIPArTT clipartt and WATT watt . Table 1 reports top-1 accuracy under synthetic corruptions and real domain shifts. We highlight two key findings. First, ClipTTA substantially improves over zero-shot CLIP, with large gains when initial accuracy is low: +20.5 pts on CIFAR-10-C, +17.4 pts on CIFAR-100-C, and +15.6 pts on Imagenet-C. In contrast, other methods perform poorly under the same conditions, which can be attributed to the increased likelihood of class collapse and pseudo-label drift when initial accuracy is low. Further analysis and finer-grained experimental evidence are presented in Sec.˜A.1 and Appendix˜B, respectively. Second, ClipTTA is the only method that consistently achieves top performance across all benchmarks. While competing methods demonstrate strengths in specific scenarios, they fall short overall. For example, ETA performs best among the TENT-style methods on domain-shift datasets but still lags by 1.4 pts on average. Similarly, CLIPArTT is most competitive on corruption benchmarks but remains 8.3 pts behind. This trend persists across both coarse- and fine-grained datasets (Fig. 3, with extended results in Sec.˜C.3). Altogether, these results establish our soft contrastive loss as the most reliable and broadly effective objective for gradient-based TTA of CLIP.
@@
Figure 3: TTA results on semantic datasets. Top-1 accuracy on coarse-grained datasets (CIFAR-10 and CIFAR-100) (a) and on $1$ fine-grained datasets (including Imagenet) (b). Comparison with gradient-based TTA methods tent ; eta ; sar ; rotta ; clipartt ; watt and alternative state-of-the-art TTA methods tpt ; tda .
@@
We further benchmark ClipTTA against two recent state-of-the-art TTA methods tailored to CLIP: TPT tpt , which adapts through text prompt tuning instead of updating normalization parameters, and TDA tda , a gradient-free approach that adjusts CLIP’s logits using cached predictions. As shown in Table 2, ClipTTA improves top-1 accuracy by an average of +19.4 pts over TPT and +15.6 pts over TDA. It achieves state-of-the-art results on nearly all domain-shift benchmarks, with the sole exception of the ImageNet-D suite, where TDA benefits from per-dataset hyperparameter tuning. However, we note that TDA lags far behind on corruption datasets, a standard benchmark in TTA, with an average gap of over 15 pts compared to ClipTTA. Figure 3 further confirms CLIPTTA’s advantage across both coarse- and fine-grained recognition tasks. While TPT and TDA are designed for single-image batches, our analysis in Sec.˜C.3 shows that ClipTTA remains competitive even in this challenging setting, thanks to its use of the CCM memory and maintains stable performance across a wide range of batch sizes.
@@ C-10-C C-100-C Imagenet-C Average VisDA-C PACS OfficeHome Imagenet-D Average
Table 3 presents results on the open-set scenario on Imagenet, where OOD detection needs to be performed alongside classification. First, we note that all closed-set methods, except ours, perform noticeably worse than zero-shot CLIP in OOD detection, highlighting CLIPTTA’s strong robustness to OOD sample contamination during adaptation. Notably, TENT and CLIPArTT suffer severe performance degradation in both classification and OOD detection, likely due to outlier interference in their pseudo-labeling process. Second, when equipped with our OCE loss, ClipTTA consistently outperforms specialized open-set TTA methods, which use heuristic OOD detection mechanisms, achieving +2.3 points AUC over UniEnt and +8.4 points AUC over SoTTA. Our soft contrastive objective reliably preserves and improves both accuracy and OOD detection, unlike these entropy-based methods, which tend to degrade CLIP’s initial performance. Additional results on other datasets are reported in Sec.˜C.3.